I don't undertand how to simplify radical expressions.I am in Algebra 2 and the problem is as follows:
Find the cubed root of (x^6/250). I was brainstorming and I know that 125*2=250 and 125 is a perfect cubed root, but I dont know if that is relevant. I would be really thankful if you could help me solve this problem. Thanks you so much!

Found 2 solutions by mananth, lwsshak3:
Answer by mananth(16946)
Answer by lwsshak3(11628)

I don't undertand how to simplify radical expressions.I am in Algebra 2 and the problem is as follows:
Find the cubed root of (x^6/250). I was brainstorming and I know that 125*2=250 and 125 is a perfect cubed root, but I dont know if that is relevant. I would be really thankful if you could help me solve this problem.
**
For this problem, assume √ is the symbol for the cube root
√(x^6/250
leave everything under the cube root sign
√(x^6/250=√(x^6/125*2)=√(x^6/5^3*2)
√(x^6/5^3*2)
multiply numerator and denominator inside by 2^2 to rationalize denominator
√(x^6*2^2/5^3*2*2^2)
√(x^6*4/5^3*2^3)
take out cube roots by dividing exponents by the index 3
x^2√4/5*2
x^2√4/10
This reads: x-square cube root of 4 divided by 10